OpenStudy (acespeedfighter):

The solution to 4.2x = 19.32 is x = ___. (3 points) 0.28 0.45 4.6 15.12

OpenStudy (agent0smith):

Divide both sides by 4.2, to get x by itself. What do you get?

OpenStudy (kittiwitti1):

\[4.2x=19.32\]\[\frac{4.2x}{4.2}=\frac{19.32}{4.2}\]\[\frac{\cancel{4.2}x}{\cancel{4.2}}=\frac{19.32}{4.2}\]\[x=\frac{19.32}{4.2}=?\]

MARC (marc_d):

solve it! @AceSpeedFighter